OpenStudy (anonymous):

someone who knows how to write a sin/cos equation from looking at the graph please reply to this.

OpenStudy (anonymous):

With this problem

OpenStudy (cggurumanjunath):

@ganeshie8 @jdoe0001 @zepdrix

OpenStudy (dumbcow):

ok lets start with general sine function \[f(x) = Asin(Bx+C)+k\] A=amplitude k= verical shift B: defined by period --> B = 2pi/period period = length of 1 cycle C : C/B is horizontal shift from graph you can identify these variables hence identify the equation

OpenStudy (jdoe0001):

the graph is shifted upwards, thus it has a "vertical shift" is also stretched some, so it has an "amplitude"

OpenStudy (anonymous):

a=2 ?

OpenStudy (jdoe0001):

and the regular period is shrunk, so it has a B component

OpenStudy (anonymous):

no a=4?

OpenStudy (jdoe0001):

amplitude looks like 2, if you look at the graph you'd count from the half-way of the wave graph because from half-up is one amplitude amount, and half-down is another

OpenStudy (anonymous):

y=2sin(2(pi/1)

OpenStudy (dumbcow):

@fanofhelp12 , A = 4

OpenStudy (jdoe0001):

ohh right... 2 above the 0 and 2 below, = 4

OpenStudy (jdoe0001):

notice the vertical shift

OpenStudy (dumbcow):

a = (max-min)/2 = 4 k = (max+min)/2 = 2

OpenStudy (anonymous):

y=4sin(2(pi)/1x) and there is no horizontal shift and the vertical shift is 2?

OpenStudy (jdoe0001):

the graph is up by 2 units, so y = 4 sin(.....) + 2

OpenStudy (anonymous):

so the period is 1 since the period ends at one..correct?

OpenStudy (dumbcow):

correct \[= 4\sin(2\pi x) +2\]

OpenStudy (anonymous):

and no horizontal so yeah y=4sin(2(pi)x)+2 THANK YOU! idk how to work opendstudy and how to give medals and such.

OpenStudy (dumbcow):

haha click the blue button "best Response"

OpenStudy (dumbcow):

unless you are on a smart phone/tablet ive heard sometimes they dont show

OpenStudy (anonymous):

Ohhhhh so i can only give one person a medal? i cant give @jdoe0001 one too? :/

OpenStudy (dumbcow):

yep new rules...give it to @jdoe0001

OpenStudy (anonymous):

its ok youve helped too:) thank you so much!

OpenStudy (anonymous):

oh i see all of your responses became best response, thanks!

OpenStudy (dumbcow):

:)